### v1

# ELO-Rated Sequence Rewards: Advancing Reinforcement Learning Models

Abstract
Reinforcement Learning (RL) is highly dependent on the meticulous design of the reward function. However, accurately assigning rewards to each state-action pair in Long-Term RL (LTRL) challenges is formidable. Consequently, RL agents are predominantly trained with expert guidance. Drawing on the principles of ordinal utility theory from economics, we propose a novel reward estimation algorithm: ELO-Rating based RL (ERRL). This approach is distinguished by two main features. Firstly, it leverages expert preferences over trajectories instead of cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward. Secondly, a new reward redistribution algorithm is introduced to mitigate training volatility in the absence of a fixed anchor reward. Our method demonstrates superior performance over several leading baselines in long-term scenarios (extending up to 5000 steps), where conventional RL algorithms falter. Furthermore, we conduct a thorough analysis of how expert preferences affect the outcomes.
keywords:
Reinforcement Learning, ELO Rating, Ordinal Reward, Reward Decomposition
00footnotetext: This paper has been accepted by DDCLS 2024.
1 Introduction
Deep Reinforcement Learning (RL) has achieved remarkable progress in various fields, including Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6]. However, real-world challenges frequently involve sparse environmental feedback. For example, a Go game may consist of approximately 200 decision states, yet there is only a signal of reward from the last state. This scarcity of feedback requires the careful design of reward functions for RL algorithms. Furthermore, the vast state spaces in many games hinder the ability of agents to learn effectively through self-exploration alone.
This situation often calls for domain experts to design the reward function [29], or use imitation learning for pre-training to guide the agent to learn effective policies [1, 30]. Nevertheless, reward functions designed by human experts face three main issues: Firstly, there may be some errors or hidden flaws in the reward function, leading RL agents to exploit these errors to obtain unreasonable rewards; Secondly, the design by human experts might cap the potential of RL agents, making the agent difficult to surpass the top human experts; Lastly, addressing large-scale problems typically requires a team of experts. These not only incurs high costs but also tends to produce policies that lack distinctive styles or characteristics due to the need for consensus of experts.
To address these issues, we took inspiration from the ordinal utility in economics, which states that people do not accurately calculate the actual value of each decision when making a decision, but obtain a decision by comparing the pros and cons of each strategy [14]. It is reasonable to assume that a RL agent may not require exact numerical rewards for effective learning. The preferences of different trajectories might help train an equally or more powerful RL agent.
Figure 1: Unlike classical RL, preference-based RL has no direct reward. Expert judgment is required when sampling trajectories from the replay buffer, and reward are obtained through expert judgment.
Preference based RL (PBRL) is the embodiment of this idea: the teacher (human expert) provides his/her the preference for the behavior of two agents (or trajectories), and then the agents use this feedback to update their policies respectively (see Figure 1). Compared with designing a reward function for each state-action pair, PBRL can not only greatly reduce the workload of human experts, but it also makes estimated reward less biased and harder to be exploited. Furthermore, it may guide the different styles of the agent through human preferences, so that humans can better understand the reasons why the agent chooses some certain strategies.
However, there are currently three problems with PBRL . First, due to the possibility of experts making mistakes, the update method of PBRL training is extremely unstable; Second, the preference judgment of experts is based on trajectories, but current RL algorithms are more suitable for the dense reward model; Third, PBRL relies solely on an expert’s preference between two trajectories, rather than leveraging the interrelationships among multiple trajectories to enhance learning.
Therefore, this paper proposes a new PBRL baseline algorithm, ELO rating based RL (ERRL). For converting the ordinal reward into cardinal reward, we adopt the ELO ranking system, which is widely used to measure player strength in zero-sum games. Particularly, we consider human experts’ preferences for trajectories as a zero-sum game, in which the preferences of different trajectories can be converted into the ELO rating for model optimization. To address the issue of sparse trajectory feedback, we utilize maximum likelihood estimation, achieving a more stable reward decomposition algorithm compared to least squares estimation. Additionally, we take some engineering optimization methods to increase the algorithm efficiency. In the experiments, the results on the selected Atari games are comparable to the state-of-the-art results in the literature. We further demonstrate the effectiveness of the proposed ERRL algorithm through the ablation experiments.
2 Preliminaries
2.1 Episodic RL with Trajectory Feedback
In classic RL settings, the research focus in Markov Decision Process (MDP) model, MDP general defined as a tuple:$\langle\mathcal{S},\mathcal{A},T,R\rangle$: where $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of environment states and action set of agent. $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ and $R(s,a)$ denote the probability of state transition to $s^{\prime}$ and the value feedback when action $a$ is taken on state $s$.
The goal of RL is to find a policy $\pi:~{}\mathcal{S}\rightarrow\mathcal{A}$ maximizing long term reward:
$\max G(s_{t},a_{t})=\max{\sum_{k=1}^{T}\gamma^{k}R(s_{t+k},a_{t+k})}$ (1)
Where $\gamma$ is discount factor.
In real-world RL problem, it is difficult for people to get the reward function $R(s,a)$ of the MDP model, we just can get the result of the game. So, in most cases, RL must face the difficulties of trajectory feedback or sparse rewards. A trajectory segment is a sequence of state-action pairs $\sigma\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$, trajectory feedback RL model general defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17], Unlike the MDP model, the environment only gives an evaluation of the entire trajectory $\mathcal{R}(\sigma)$ at the end of the game.
$\mathcal{R}(\sigma)=\sum_{(s_{i},a_{i})\in\sigma}R(s_{i},a_{i})$ (2)
In order to improve training efficiency, it is necessary to redistribute the entire reward to each state-action pair [15, 17], convert the trajectory feedback RL problem into a standard form of RL problem. Constructing a proxy reward function $\hat{r}$ to estimate reward on state-action pair like MDP model. In practice, we optimize the proxy reward function using the following loss function.
$\mathcal{L}(\hat{r})=\mathbb{E}_{\sigma\in\mathcal{D}}\left[\left(\mathcal{R}(\sigma)-{\sum}_{(s_{i},a_{i})\in\sigma}\hat{r}(s_{i},a_{i})\right)\right]$ (3)
2.2 Ordinal Utility Theory and Cardinal Utility Theory
Cardinal utility theory is a popular concept in microeconomics during the 19th and early 20th centuries [18]. The basic idea is that utility can be measured and summed up. The base unit of utility is called the utility unit. Utility unit can be expressed numerical e.g. $(1,2,3\dots)$, just as the unit of length can be expressed in meters. The mainstream RL reward function $R(s,a)$ is similar to the definition of Cardinality Utility,
Ordinal utility theory was proposed by Hicks and Allen in 1934 [14], and the main point is that: utility as a psychological phenomenon cannot be measured because it is impossible to find a reasonable unit for measurement utility; They believe that what consumers do in the market is not to measure the utility of goods, but to rank among different goods.
Similarly, in some RL problems, it is very difficult to get an accurate Cardinality Reward, but it is easier to obtain the preferences on different trajectories. For example, there are generally 80$\sim$100 decision points in chess, it is difficult to estimate the numerical value of particular states even professional chess players, but the final match result of chess can be judged by even beginner.
2.3 Elo Rating System
The ELO rating system  [21] is an evaluation method created by Arpad Elo to measure the level of players in various zero-sum game competitions. It is a recognized authoritative method for evaluating players’ abilities widely in competitive competitions, such as chess, Go, football, basketball, e-sports and other competitive competitions. Each player has a hidden rating score, which represents the player’s average competitive level. The higher the score, the stronger the player’s ability. However, due to the randomness of the game and the status of the players, there often exists a gap between the on-the-spot scores and the hidden scores. At the ELO rating model, the hidden score needs to add a fluctuation error, to represent the real ability of the on-the-spot performance. The purpose of the ELO rating is to estimate the true abilities (hidden score) of the players through the win-lose relationship as accurately as possible . The final ELO rating model is as follows:
Random variable $\xi$ obey Logistic distribution, denote the gap between on-the-spot scores and the hidden scores:
$P(\xi\leq x)=F(x)=\frac{1}{1+e^{-x}}$ (4)
$R_{A}$ and $R_{B}$ denote the current hidden score of Player A and Player B, let $D=R_{A}-R_{B}$, At this time, the probability of player A defeat player B is:
$P(D)$ $=1/2+F(x)\big{|}_{0}^{\frac{D}{2}}$ (5)
$=\frac{1}{1+e^{\frac{D}{2}}}$
In order to facilitate calculation and to show the ability gap between players intuitively, ELO modifies the variance of the Logistic distribution and set the initial score at 1500.
Assume the win rate of player A is $E(\sigma^{A},\sigma^{B})$:
$E(\sigma^{A},\sigma^{B})=\frac{1}{1+10^{(R_{B}-R_{A})/400}}$ (6)
$\mu$ denote result function, if $\mu(A)=1$ denote player A defeated player B, $\mu(A)=0$ denote player B defeated player A, Then player A’s ELO rating is updated to:
$R_{A}^{\prime}=R_{A}+K(\mu(A)-E(\sigma^{A},\sigma^{B}))$ (7)
2.4 Performance Based RL (PBRL)
PBRL [16] is a typical algorithm using ordinal utility in RL, which divides the entire trajectory into many sub-trajectories with fixed length $k$, and using $\sigma^{A}\succ\sigma^{B}$ indicates that human experts prefer sub-trajectory $\sigma^{A}$ to sub-trajectory $\sigma^{B}$.
If human experts judge $\sigma^{A}\succ\sigma^{B}$, then there is a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
So it is reasonable to assume that under the proxy reward function $\hat{r}$, the probability of an experts making $\sigma^{A}\succ\sigma^{B}$ is:
$\hat{P}(\sigma^{A}\succ\sigma^{B})=\frac{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})}{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})+\exp\sum\limits_{i=1}^{i=k_{B}}\hat{r}(s^{B}_{i},a^{B}_{i})}$ (8)
Under this assumption, store expert’s preference results $\left(\sigma^{A},\sigma^{B}\,\mu\right)$ into the database $\mathcal{D}$, where $\mu$ is the result function, if $\mu(A)=1$ means $\sigma^{A}\succ\sigma^{B}$, $\mu(A)=0$ means $\sigma^{B}\succ\sigma^{A}$, according to the preference, use the cross entropy loss to optimize the proxy reward function $\hat{r}$ and train reinforcement learning through the proxy reward function $\hat{r}$:
$\mathcal{L}(\hat{r})=-\sum_{(\sigma^{A},\sigma^{B}\,\mu)\in\mathcal{D}}\left(\begin{aligned} \mu(A)\log\hat{P}(\sigma^{A}\succ\sigma^{B})+\\ \mu(B)\log\hat{P}(\sigma^{B}\succ\sigma^{A})\end{aligned}\right)$ (9)
3 Methodology
In this section, we introduce our new method, ELO rating based RL. It abandons the numerical reward, sorts the trajectories through ELO rating, and uses a simplified version of the return decomposition method to assign ELO rating to each state-action pair.
3.1 Trail Reward Estimation Via Elo Rating
Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose that in trajectory $\sigma^{A}$ achieves better results than trajectory $\sigma^{B}$ formulating as $\sigma^{A}\succ\sigma^{B}$ (ie. in Chess, player A wins Player B). Denote $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$ as the reward obtained by this trajectory $\sigma^{X}$, where $\overline{\mathcal{R}(\sigma^{X})}$ represents the expected reward of using the trajectory $\sigma^{X}$, $\epsilon$ denote the reward fluctuation. In general, it is very difficult to estimate $\overline{\mathcal{R}(\sigma^{X})}$, either need to consume huge computing resources to fully explore the entire solution space; or hire human experts to design appropriate intermediate reward to guide the agent to explore more efficiently.
Although it is difficult to give the appropriate reward for each trajectory $\mathcal{R}(\sigma^{X})$, it is much simpler to assess the pros and cons of each trajectory. At this time, it is better to use the idea of ordinal utility, using $\mathcal{U}(\sigma)$ to represent the ELO rating of the trajectory $\sigma$, and use $\mathcal{U}(\sigma)$ as the reward of the entire trajectory to guide RL training. When $\sigma^{A}\succ\sigma^{B}$, it can be regarded as player $\sigma^{A}$ defeated player $\sigma^{B}$ ($S_{A}=1$); if the expert cannot judge the pros and cons of the two trajectories, it can be described a draw, $\sigma^{A}\approx\sigma^{B}$ ($\approx$ indicates that the experts believe that the results of the two trajectories are similar), can be regarded as a draw between the two sides ($S_{A}=0.5$) . If $\sigma^{A}\succ\sigma^{B}$ the event $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$ will occur with high probability, and the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ can be updated to:
$E(\sigma^{A},\sigma^{B})=\frac{1}{1+10^{(\hat{\mathcal{U}}(\sigma^{B})-\hat{\mathcal{U}}(\sigma^{A}))/\eta}}$ (10)
$\hat{\mathcal{U}}^{\prime}(\sigma^{A})=\hat{\mathcal{U}}(\sigma^{A})+K(S_{A}-E(\sigma^{A},\sigma^{B}))$ (11)
Where $\eta$ and $K$ are the parameters to adjust the training fluctuation.
3.2 Indirect Redistribution of Long-Term Reward
$\hat{\mathcal{U}}(\sigma)$ is estimating on the entire trajectory, which should be the sum of the ELO rating for all state-action pairs:
$\hat{\mathcal{U}}(\sigma^{X})=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (12)
In value based RL, the value function itself implies a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, there is no need to redefine a proxy reward function, We use deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate long-term ELO rating on the state-action pair $(s,a)$. The score for the entire trajectory $\sigma^{X}$ is denote as $\hat{\mathcal{U}}(\sigma^{X})$:
$\hat{\mathcal{U}}(\sigma^{X})$ $=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (13)
$=\sum_{t=1}^{k_{X}}\left[\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})-\gamma\hat{G}_{\text{elo}}(s^{X}_{t+1},a^{X}_{t+1})\right]$
The long-term reward for each state-action pair is updated as (proved in Appendix A):
$\hat{G}^{\prime}_{\text{elo}}(s^{X}_{t},a^{X}_{t})=\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})+\frac{K(S_{X}-E_{X})}{k_{X}}$ (14)
3.3 Combining RL and Elo Rating
At the code implementation level, our algorithm is compatible with mainstream RL algorithms. Our algorithm migrates perfectly to any RL algorithm with only a modification to the reward function. In Algorithm 1, we integrate actor-critic RL with Elo rating.
Algorithm 1 AC RL with Elo Rating
1:Initialize $\mathcal{D}\leftarrow\varnothing$
Predictably, the preferences of experts will have a large impact on the results of ordinal-based RL . In addition, the $K$ and $\eta$ parameters of ELO rating will also have a great impact on the results theoretically. We will discuss this problem in detail in Section 4 .
4 Experiments
In this section, we conduct experiments on a suite of Atari benchmark tasks to investigate the empirical performance of our proposed method. We compare our method with several baseline algorithms in the literature, and conduct an ablation study.
4.1 Performance in Atari Benchmark with episodic preference
4.1.1 Setups
To evaluate the behavior of ERRL, we compared its performance with several based algorithms across four Atari tasks. In ordinal based RL, all trajectories record some key events for expert judgment instead of returning a numerical reward. These tasks are long-term feedback problems (up to 5000 steps). We use the open source RL framework RLLib [24] as the baseline, which is an efficient parallel RL framework and adopts the standard preprocessing of Deepmind [7]. All results have been fine-tuned, and detailed settings can be found in Appendix B.
Since our experiment is based on ordinal RL, how to generate the ordinal model is crucial to the learning results. Here we take Pong as an example to illustrate the ordinal production process.
Pong is a game that simulates table tennis (ping-pong). In normal mode, if you shot the ball out of the opponent’s baseline, you will get 1 point; if you don’t catch the opponent’s shot, you will deduct 1 point. The player who reaches 21 points first wins the game. The final trajectory reward is $\mathcal{R}(\sigma)$, its own score minus the opponent’s score. And our evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
$k$ indicates the length of the trajectory. In order to avoid the trajectory being too long, the game automatically terminates after the step reaches 5000.
The rationale behind this design is twofold: if defeating your opponent directly is not feasible, the strategy shifts towards dragging your opponent into extended play, aiming to hit as many of the opponent’s balls as possible. Conversely, when you can defeat your opponent, the approach is to leverage any means necessary to secure the win. So there are no more restrictions. For more game settings please refer to Appendix B.
4.1.2 Performance in Atari Benchmark
Figure 2: Learning curves on a suite of Atari benchmark tasks with trajectory feedback. The shaded region indicates 95% confidence interval. Due to the trajectory length is indeterminate, we set up an evaluation point every 60 trajectories, about $3\times 10^{4}\sim 3\times 10^{5}$ time-steps.
A comparison of overall performance, as shown in Figure 2, shows that ERRL generally outperforms the baseline algorithms on these several problems. Our method can achieve higher sample efficiency and generate better policies after convergence. Because the returns are sparse, it is difficult for the original PPO to learn useful information. As for the previous reward decomposition function, it is mainly aimed at the decomposition of numerical reward. The reward obtained directly from the conversion of ordinal reward will increase the fluctuation of training. When the agent employs PPO for training with the reward conversion based on ELO Rating, it can learn an effective policy. This is because the ELO rating tends to amplify the performance gap between trajectories. In the long-term RL problems, ELO ratings are more conducive to highlighting the pros and cons of trajectories than the compact real rewards. So that the agent can find useful information.
4.2 Ablation Study
We study the value of $\eta$ in the ERRL algorithm under the condition of a fixed learning rate of $1\times 10^{-4}$. For the value of $K$, we refer to the original design of ELO rating to fix $K=0.04\times\eta$.
As described in Section 3, $\eta$ controls the fluctuation range of the ELO rating, smaller $\eta$ represents greater randomness of the zero-sum game, larger training fluctuations, but training faster; larger $\eta$ represents less randomness, and reduced training fluctuations, but training slower. So we need a suitable $\eta$ to balance training speed and fluctuation.
Figure 3: Learning curves of ERRL in Pong with different choices of hyper-parameter $\eta$. The curves with $\eta=0.01$ correspond to the default implementation.
We perform an ablation study on the hyperparameter $\eta$ that represents the range of fluctuations in the ELO rating. Considering that the time steps of the game will change with the progress of the game, we decided to relating $\eta$ with the time-steps of the trajectory, that is, the $\eta$ actually used in the ELO Rating calculation is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$. For the convenience of writing, we will use $\eta$ to represent $\eta_{\text{our}}$ in the following text if there is no special emphasis.
We evaluate ERRL with a set of choices of hyperparameter $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$. The experimental results shown in Figure 3 show that this hyperparameter has a great impact on RL training. A larger $\eta$ will cause the training fluctuation to be too large and directly collapse, while a smaller $\eta$ will cause the training to be too slow and easily lead to over fitting. An acceptable range of $\eta$ is in interval $[0.05,0.005]$.
From the time-steps of the entire trajectory, regardless of the size of $\eta$, the final time-steps converge to 2100$\sim$2400. It can be seen that under the same preference, if the training can converge, the choice of $\eta$ does not change the final strategies, and we will discuss the relationship between strategy selection and expert preference in Appendix C.
5 Related Work
RL without Environmental Reward. While RL typically relies on environmental reward, several studies have explored RL without direct environmental rewards. This research can be broadly categorized into two approaches. The first approach focuses on capturing essential information within trajectories and using these differences to define learning rewards Eysenbach et al. [28] proposed a pre-training method to train the discriminator to identify which agent generated the trajectory. the agent always hopes that the discriminator can easily identify itself, so the agent always tends to find a unique state. Hu et al. [26] uses multidimensional features to determine the Pareto frontier. Only the agents at the Pareto frontier will perform the next training. This approach enables game designers to create a variety of agents with distinct styles, but this type of method is generally only suitable for increasing the diversity of agents. Another idea is to use human embeddings in RL, such as PBRL [16] introduced in Section 2.4. In addition, Kimin et al. [13] proposed the method of PEBBLE, which has demonstrated exceptional outcomes. .
Sparse Rewards RL. Sparse rewards significantly hinder the application of RL [11, 8], with numerous studies addressing this challenge. A straightforward solution is the integration of human experience [1, 4]. As long as the human benefit is designed perfectly enough, this approach can lead to successful outcomes. However, relying heavily on human expertise can be expensive and may inadvertently limit the capabilities of the RL agent [1, 4, 5]. Another approach is reward decomposition, which decomposes the total reward into individual state-action pairs [15, 20]. usually this method requires reward shaping to achieve good results. Marcin et al. [10] also achieves good results by reusing historical information, but this only works for tasks with simple and clear states.
6 Conclusion and Future Work
We propose a new preference-based RL algorithm ERRL, which can achieve good performance in the 5000-step long-tern trajectory feedback RL problem. ERRL abandons the classical numerical reward, and estimates the ordinal reward of each trajectory through the expert’s preference and the ELO rating method. In addition, a new reward decomposition method is adopted to reduce the fluctuation of training in ERRL setting. Experiments show that ERRL outperforms previous methods in solution quality. In future work, we will explore the theoretical basis for combining ELO rating with RL, conduct more detailed experiments, discuss the effectiveness of our algorithm when it is difficult for experts to make judgments, and explore the feasibility of ELO rating in the remaining areas. Such as reward shaping  [27], games with hard-to-determine payoffs (ie. Texas hold’em and Hanabi) [25], or research on the diversity of RL agents [26, 28].
Appendix A
$\hat{\mathcal{U}}(\sigma^{A})$ is the ELO rating estimate of the entire trajectory $\sigma^{A}$, in trajectory $\sigma^{A}$, let the long-term ELO rating $\hat{U}(s_{t}^{X},a_{t}^{X})$ for each state-action pair update as $\delta_{t}^{X}$,
$\hat{U}^{\prime}(s_{t}^{A},a_{t}^{A})=\hat{U}(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}$ (15)
In theory, any update can be regarded as an effective update range as long as it is randomly sampled so that the $\delta_{t}^{X}$ distribution satisfies the $K(S_{X}-E_{X})=\delta_{t}^{X}$ constraint. But this obviously leads to too large variance, which is not suitable for iterative calculation. So we might as well consider $\delta_{t}^{A}$ as the assumption of the ELO rating on the player’s ability to fluctuate, and assume that the estimation error for each state-action pair is independent and uniformly obeys the Logistic distribution:
$\delta_{t}^{A}\sim\text{Logistic}(x)=\frac{e^{-x}}{(1+e^{-x})^{2}}$ (16)
According to maximum likelihood estimation, the above problem can be transformed into an optimization problem. The current formula  17 takes the maximum value of $x_{t}$ as the most likely update range:
$\begin{split}\max\prod_{t=1}^{k_{A}}\frac{e^{-\delta_{t}^{A}}}{(1+e^{-\delta_{t}^{A}})^{2}}\\ \text{s.t.}~{}~{}~{}~{}(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (17)
Taking the logarithm of the original formula 17, there are:
$\max-\sum_{t=1}^{k_{A}}\left(\delta_{t}^{A}+2\ln(1+e^{-\delta_{t}^{A}})\right)$ (18)
Bringing into $\sum_{t=1}^{k_{A}}=(S_{A}-E_{A})$
$~{}~{}~{}\sum_{t=1}^{k_{A}}\left(\ln(1+e^{-\delta_{t}^{A}})\right)$ (19)
$=\sum_{t=1}^{k_{A}}\left(\frac{\ln(1+e^{\delta_{t}^{A}})}{e^{\delta_{t}^{A}}}\right)$
Finally:
$\begin{split}\min\sum_{t=1}^{k_{A}}\ln(1+e^{\delta^{A}_{t}})\\ \text{s.t.}~{}~{}~{}~{}K(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (20)
$\begin{split}\min\sum_{t=1}^{k_{A}}\ln(1+e^{\delta^{A}_{t}})\\ \text{s.t.}~{}~{}~{}~{}K(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (21)
It is easy to know that when $\delta_{t}^{A}=\frac{K(S_{A}-E_{A})}{k_{A}}$, the cost function obtains the minimum value, so the new model is updated as:
$\hat{G}_{\text{elo}}^{\prime}(s_{t}^{A},a_{t}^{A})=\gamma\hat{G}_{\text{elo}}(s_{t+1}^{A},a_{t+1}^{A})+U(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}\$ (22)
Bringing into $U(s_{t}^{A},a_{t}^{A})=\hat{G}_{\text{elo}}(s_{t}^{A},a_{t}^{A})-\gamma\hat{G}_{\text{elo}}(s_{t+1}^{A},a_{t+1}^{A})$,
$\hat{G}_{\text{elo}}^{\prime}(s_{t}^{A},a_{t}^{A})=\hat{G}_{\text{elo}}(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}$ (23)
Appendix B
Hyper-Parameter Configuration For Atari
In terms of hardware configuration, the CPU is AMD thread-rippers 3990WX with 32 cores and 64 threads, the graphics card is Nvidia RTX 3080Ti, and the memory is 128GB . During training, 60 workers are used to collect data at the same time to improve the experimental efficiency. The training method of agent is PPO . We refer to the default settings of PPO [24, 22], the main hyperparameter settings are shown in the figure 1.
Table 1: Hyper-Parameter
Hyper-Parameter Default Configuration
discount factor $\gamma$ 0.99
activation ReLU
stacked frames in agent observation 4
state mode memory
actor agent num 60
network architecture PPO
optimizer Adam
learning rate $1\cdot 10^{-4}$
gradient steps per environment step 30000$\sim$300000
ELO $\eta$ 0.01
ELO $K$ $0.04\times\eta$
Our open source main experimental code and details can be found in https://github.com/Zealoter/ERRL, as our experiments are still ongoing, we will gradually improve our code and experimental details.
Expert Preferences
Pong
In Pong game, you control the right paddle, you compete against the left paddle controlled by the computer. Unlike reality ping-pong, the upper and lower sides of the table have boundaries, regardless of the ball going out of bounds. Evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
SpaceInvaders
In this game, your goal is to fire laser cannons at space invaders to destroy them before they reach Earth. The game is over when you lose all your lives under enemy’s fire, or when they reach Earth.Since the aliens will gradually approach the earth unless they are eliminated. Obviously, the score is not as important as the time in this game. This is because the agent may give up killing the front row in order to kill the high-scoring aliens in the back row. So the evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $K_{A}>K_{B}$ or $K_{A}<K_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
BeamRider
In this game, you control a space-ship that travels forward at a constant speed. You can only steer it sideways between discrete positions. However, at the end of the level only torpedoes can defeat the boss, and the number of torpedoes is limited, so the preferences of this game are difficult to design. From the player’s point of view it’s always beneficial to live longer, so we combined time-steps length and score to design BeamRider’s preferences. The evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})+K_{A}>\mathcal{R}(\sigma^{B})+K_{B}$ or $\mathcal{R}(\sigma^{A})+K_{A}<\mathcal{R}(\sigma^{B})+K_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Appendix C
Here we discuss several settings of expert preference using Pong and explore the impact of different expert preferences on training.
Preference 1: normal mode
This mode is the same as the previous.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 2: reward mode
Figure 4: It can be seen that the preference has a great influence on ERRL, and changing the preference may change the final policy of the agent.
In this mode, preference comes only from reward.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 3: ball control mode
In this mode, trajectories with more time-steps are always better in the same reward. That’s why it’s called ball control mode.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 4: aggressive mode
We want the agent to beat the opponent as quickly as possible, so when agent is able to beat the opponent, less time-steps is the better.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then see the third item, if $\mathcal{R}(\sigma^{A})<0$ and $\mathcal{R}(\sigma^{B})<0$ then see the fourth item, $\sigma^{A}\approx\sigma^{B}$.
  3. 3.
    If $k_{A}<k_{B}$ or $k_{A}>k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
  4. 4.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 5: sudden death mode
In this mode, if $\mathcal{R}(\sigma^{X})\leq-5$, game over directly, and the rest is the same as preference 1.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
It can be seen that inappropriate preferences may fail to train useful agents. In the case of ball control mode, the agent gets a policy that can play more time-steps in the game. However, the rest of the models did not reach satisfactory conclusions, so we will explore the relationship between expert preference and final strategy in future research.

### v2

@@
 Abstract
-Reinforcement Learning (RL) is highly dependent on the meticulous design of the reward function.
-However, accurately assigning rewards to each state-action pair in Long-Term RL (LTRL) challenges is formidable.
-Consequently, RL agents are predominantly trained with expert guidance.
-Drawing on the principles of ordinal utility theory from economics, we propose a novel reward estimation algorithm: ELO-Rating based RL (ERRL).
-This approach is distinguished by two main features.
-Firstly, it leverages expert preferences over trajectories instead of cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward.
-Secondly, a new reward redistribution algorithm is introduced to mitigate training volatility in the absence of a fixed anchor reward.
-Our method demonstrates superior performance over several leading baselines in long-term scenarios (extending up to 5000 steps), where conventional RL algorithms falter.
-Furthermore, we conduct a thorough analysis of how expert preferences affect the outcomes.
+Reinforcement Learning (RL) heavily relies on the careful design of the reward function.
+However, accurately assigning rewards to each state-action pair in Long-Term Reinforcement Learning (LTRL) tasks remains a significant challenge.
+As a result, RL agents are often trained under expert guidance.
+Inspired by the ordinal utility theory in economics, we propose a novel reward estimation algorithm: ELO-Rating based Reinforcement Learning (ERRL).
+This approach features two key contributions.
+First, it uses expert preferences over trajectories rather than cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward.
+Second, a new reward redistribution algorithm is introduced to alleviate training instability in the absence of a fixed anchor reward.
+In long-term scenarios (up to 5000 steps), where traditional RL algorithms struggle, our method outperforms several state-of-the-art baselines.
+Additionally, we conduct a comprehensive analysis of how expert preferences influence the results.
 keywords:
 Reinforcement Learning, ELO Rating, Ordinal Reward, Reward Decomposition
-00footnotetext: This paper has been accepted by DDCLS 2024.
 1 Introduction
-Deep Reinforcement Learning (RL) has achieved remarkable progress in various fields, including Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6].
-However, real-world challenges frequently involve sparse environmental feedback.
-For example, a Go game may consist of approximately 200 decision states, yet there is only a signal of reward from the last state.
-This scarcity of feedback requires the careful design of reward functions for RL algorithms.
-Furthermore, the vast state spaces in many games hinder the ability of agents to learn effectively through self-exploration alone.
-This situation often calls for domain experts to design the reward function [29], or use imitation learning for pre-training to guide the agent to learn effective policies [1, 30].
-Nevertheless, reward functions designed by human experts face three main issues: Firstly, there may be some errors or hidden flaws in the reward function, leading RL agents to exploit these errors to obtain unreasonable rewards; Secondly, the design by human experts might cap the potential of RL agents, making the agent difficult to surpass the top human experts; Lastly, addressing large-scale problems typically requires a team of experts.
-These not only incurs high costs but also tends to produce policies that lack distinctive styles or characteristics due to the need for consensus of experts.
-To address these issues, we took inspiration from the ordinal utility in economics, which states that people do not accurately calculate the actual value of each decision when making a decision, but obtain a decision by comparing the pros and cons of each strategy [14].
-It is reasonable to assume that a RL agent may not require exact numerical rewards for effective learning.
-The preferences of different trajectories might help train an equally or more powerful RL agent.
+Deep Reinforcement Learning (RL) has witnessed significant advancements across diverse domains, such as Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6].
+However, real-world applications often encounter sparse environmental feedback.
+For instance, in a game of Go, which may involve around $N\approx 200$ decision states, the reward signal is solely available at the final state.
+This sparsity necessitates meticulous design of reward functions for RL algorithms.
+Moreover, the expansive state spaces in many games impede agents from effectively learning through self-exploration.
+In response, it is common practice to either engage domain experts in reward function design [29] or employ imitation learning for pre-training to guide agents in learning optimal policies [1, 30].
+Nevertheless, reward functions crafted by human experts are confronted with three primary challenges: First, potential errors or latent flaws within the reward function can be exploited by RL agents to obtain spurious rewards; Second, expert-designed functions may impose an upper bound on the performance of RL agents, hindering them from surpassing human expertise; Third, tackling large-scale problems usually demands a team of experts, which not only escalates costs but also often results in policies lacking uniqueness due to the requirement for expert consensus.
+To address these challenges, we draw inspiration from the ordinal utility theory in economics, which posits that decision-makers do not compute the exact numerical value of each choice, but rather derive decisions by comparing the relative merits of strategies [14].
+This motivates the hypothesis that RL agents may not require precise cardinal rewards for effective learning—instead, preferences over trajectories could suffice to train agents with comparable or superior performance.
 Figure 1: Unlike classical RL, preference-based RL has no direct reward. Expert judgment is required when sampling trajectories from the replay buffer, and reward are obtained through expert judgment.
-Preference based RL (PBRL) is the embodiment of this idea: the teacher (human expert) provides his/her the preference for the behavior of two agents (or trajectories), and then the agents use this feedback to update their policies respectively (see Figure 1).
-Compared with designing a reward function for each state-action pair, PBRL can not only greatly reduce the workload of human experts, but it also makes estimated reward less biased and harder to be exploited.
-Furthermore, it may guide the different styles of the agent through human preferences, so that humans can better understand the reasons why the agent chooses some certain strategies.
-However, there are currently three problems with PBRL .
-First, due to the possibility of experts making mistakes, the update method of PBRL training is extremely unstable; Second, the preference judgment of experts is based on trajectories, but current RL algorithms are more suitable for the dense reward model; Third, PBRL relies solely on an expert’s preference between two trajectories, rather than leveraging the interrelationships among multiple trajectories to enhance learning.
-Therefore, this paper proposes a new PBRL baseline algorithm, ELO rating based RL (ERRL).
-For converting the ordinal reward into cardinal reward, we adopt the ELO ranking system, which is widely used to measure player strength in zero-sum games.
-Particularly, we consider human experts’ preferences for trajectories as a zero-sum game, in which the preferences of different trajectories can be converted into the ELO rating for model optimization.
-To address the issue of sparse trajectory feedback, we utilize maximum likelihood estimation, achieving a more stable reward decomposition algorithm compared to least squares estimation.
-Additionally, we take some engineering optimization methods to increase the algorithm efficiency.
-In the experiments, the results on the selected Atari games are comparable to the state-of-the-art results in the literature.
-We further demonstrate the effectiveness of the proposed ERRL algorithm through the ablation experiments.
+Preference-based RL (PBRL) operationalizes this concept: a teacher (human expert) provides preferences over the behaviors of two agents (or trajectories), which the agents then use to update their policies (see Figure 1).
+Compared to designing a reward function for each state-action pair, PBRL not only reduces the workload for human experts but also yields estimated rewards that are less biased and less susceptible to exploitation.
+Moreover, human preferences can guide agents to exhibit diverse styles, enabling better interpretability of why certain strategies are chosen.
+However, current PBRL approaches face three key limitations: First, training instability arises from potential expert errors in preference judgments; Second, while experts provide trajectory-level preferences, most RL algorithms are optimized for dense reward models; Third, PBRL typically relies on pairwise trajectory preferences rather than leveraging inter-trajectory relationships to enhance learning.
+To address these, we propose a novel PBRL baseline algorithm, ELO Rating-based RL (ERRL).
+For converting ordinal preferences to cardinal rewards, we adopt the ELO ranking system, widely used to measure player strength in zero-sum games.
+Specifically, we model human trajectory preferences as a zero-sum game, transforming preference signals into ELO ratings for model optimization.
+To handle sparse trajectory feedback, we employ maximum likelihood estimation (MLE), which yields a more stable reward decomposition than least squares methods.
+We also introduce engineering optimizations to enhance algorithm efficiency.
+Experiments on selected Atari games demonstrate that ERRL achieves performance comparable to state-of-the-art baselines.
+Ablation studies further validate the effectiveness of our approach.
+The code is available on GitHub.
 2 Preliminaries
 2.1 Episodic RL with Trajectory Feedback
-In classic RL settings, the research focus in Markov Decision Process (MDP) model, MDP general defined as a tuple:$\langle\mathcal{S},\mathcal{A},T,R\rangle$: where $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of environment states and action set of agent.
-$T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ and $R(s,a)$ denote the probability of state transition to $s^{\prime}$ and the value feedback when action $a$ is taken on state $s$.
+In classic RL settings, the research focuses on the Markov Decision Process (MDP) model, generally defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,R\rangle$.
+Here, $\mathcal{S}$ and $\mathcal{A}$ denote the environment state space and agent action set, respectively.
+$T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ represents the probability of transitioning to state $s^{\prime}$ when action $a$ is taken in state $s$, while $R(s,a)$ denotes the value feedback for this state-action pair.
 The goal of RL is to find a policy $\pi:~{}\mathcal{S}\rightarrow\mathcal{A}$ maximizing long term reward:
@@
 Where $\gamma$ is discount factor.
-In real-world RL problem, it is difficult for people to get the reward function $R(s,a)$ of the MDP model, we just can get the result of the game.
-So, in most cases, RL must face the difficulties of trajectory feedback or sparse rewards.
-A trajectory segment is a sequence of state-action pairs $\sigma\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$, trajectory feedback RL model general defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17], Unlike the MDP model, the environment only gives an evaluation of the entire trajectory $\mathcal{R}(\sigma)$ at the end of the game.
+In real-world RL scenarios, specifying the reward function $R(s,a)$ of the MDP model is often impractical; instead, only terminal outcomes of trajectories are observable.
+Consequently, RL agents frequently encounter challenges related to sparse or delayed feedback.
+A trajectory segment is defined as a sequence of state-action pairs $\sigma=\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$.
+In response, trajectory-feedback RL models are typically formalized as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17].
+Unlike standard MDPs, the environment provides only a cumulative evaluation $\mathcal{R}(\sigma)$ over the entire trajectory $\sigma$ upon termination.
 $\mathcal{R}(\sigma)=\sum_{(s_{i},a_{i})\in\sigma}R(s_{i},a_{i})$ (2)
-In order to improve training efficiency, it is necessary to redistribute the entire reward to each state-action pair [15, 17], convert the trajectory feedback RL problem into a standard form of RL problem.
-Constructing a proxy reward function $\hat{r}$ to estimate reward on state-action pair like MDP model.
-In practice, we optimize the proxy reward function using the following loss function.
+To enhance training efficiency, it is essential to redistribute the cumulative trajectory reward to individual state-action pairs [15, 17], thereby transforming the trajectory-feedback RL problem into a standard MDP framework.
+This involves constructing a proxy reward function $\hat{r}(s,a)$ to approximate the state-action value, analogous to the reward function in traditional MDPs.
+In practice, we optimize this proxy reward function using the following loss function:
 $\mathcal{L}(\hat{r})=\mathbb{E}_{\sigma\in\mathcal{D}}\left[\left(\mathcal{R}(\sigma)-{\sum}_{(s_{i},a_{i})\in\sigma}\hat{r}(s_{i},a_{i})\right)\right]$ (3)
 2.2 Ordinal Utility Theory and Cardinal Utility Theory
-Cardinal utility theory is a popular concept in microeconomics during the 19th and early 20th centuries [18].
-The basic idea is that utility can be measured and summed up.
-The base unit of utility is called the utility unit.
-Utility unit can be expressed numerical e.g.
-$(1,2,3\dots)$, just as the unit of length can be expressed in meters.
-The mainstream RL reward function $R(s,a)$ is similar to the definition of Cardinality Utility,
-Ordinal utility theory was proposed by Hicks and Allen in 1934 [14], and the main point is that: utility as a psychological phenomenon cannot be measured because it is impossible to find a reasonable unit for measurement utility; They believe that what consumers do in the market is not to measure the utility of goods, but to rank among different goods.
-Similarly, in some RL problems, it is very difficult to get an accurate Cardinality Reward, but it is easier to obtain the preferences on different trajectories.
-For example, there are generally 80$\sim$100 decision points in chess, it is difficult to estimate the numerical value of particular states even professional chess players, but the final match result of chess can be judged by even beginner.
+Cardinal utility theory is a prominent concept in microeconomics during the 19th and early 20th centuries [18].
+Its core idea is that utility can be measured and summed, with the basic unit of utility referred to as the utility unit.
+Utility units can be expressed numerically (e.g., $(1,2,3,\dots)$), analogous to how length is measured in meters.
+The mainstream RL reward function $R(s,a)$ mirrors the definition of cardinal utility.
+Ordinal utility theory, proposed by Hicks and Allen in 1934 [14], posits that utility as a psychological phenomenon cannot be measured, as no reasonable unit for quantifying utility exists.
+They argue that consumers in the market do not measure the utility of goods but rather rank different goods.
+Similarly, in certain RL problems, obtaining an accurate cardinal reward is highly challenging, whereas acquiring preferences over different trajectories is more feasible.
+For instance, chess typically involves 80–100 decision points, making it difficult even for professional chess players to estimate the numerical value of specific states.
+However, even beginners can judge the final match result.
 2.3 Elo Rating System
-The ELO rating system  [21] is an evaluation method created by Arpad Elo to measure the level of players in various zero-sum game competitions.
-It is a recognized authoritative method for evaluating players’ abilities widely in competitive competitions, such as chess, Go, football, basketball, e-sports and other competitive competitions.
-Each player has a hidden rating score, which represents the player’s average competitive level.
-The higher the score, the stronger the player’s ability.
-However, due to the randomness of the game and the status of the players, there often exists a gap between the on-the-spot scores and the hidden scores.
-At the ELO rating model, the hidden score needs to add a fluctuation error, to represent the real ability of the on-the-spot performance.
-The purpose of the ELO rating is to estimate the true abilities (hidden score) of the players through the win-lose relationship as accurately as possible .
-The final ELO rating model is as follows:
-Random variable $\xi$ obey Logistic distribution, denote the gap between on-the-spot scores and the hidden scores:
+The ELO rating system [21] is an evaluation framework developed by Arpad Elo to measure player performance in zero-sum games.
+Widely adopted in competitive domains such as chess, Go, football, basketball, and esports, it serves as a standard for assessing player skill levels.
+Each player is assigned a latent rating score representing their average competitive ability: higher scores indicate stronger performance.
+Due to game randomness and player variability, observed outcomes may deviate from latent ratings.
+In the ELO model, this discrepancy is modeled as a fluctuating error term to reconcile actual performance with true ability.
+The core objective of ELO rating is to estimate player abilities (latent scores) as accurately as possible from win-loss outcomes.
+The formal ELO rating model is defined as follows:
+Let the random variable $\xi$ follow a Logistic distribution, representing the gap between observed and latent scores:
 $P(\xi\leq x)=F(x)=\frac{1}{1+e^{-x}}$ (4)
-$R_{A}$ and $R_{B}$ denote the current hidden score of Player A and Player B, let $D=R_{A}-R_{B}$, At this time, the probability of player A defeat player B is:
-$P(D)$ $=1/2+F(x)\big{|}_{0}^{\frac{D}{2}}$ (5)
-$=\frac{1}{1+e^{\frac{D}{2}}}$
-In order to facilitate calculation and to show the ability gap between players intuitively, ELO modifies the variance of the Logistic distribution and set the initial score at 1500.
+Denote $R_{A}$ and $R_{B}$ as the current latent scores of Players A and B, respectively, and let $D=R_{A}-R_{B}$. The probability that Player A defeats Player B is:
+[…]
+$=\frac{1}{1+e^{-\frac{D}{2}}}$
+To facilitate computation and intuitively reflect ability differences, ELO modifies the Logistic distribution’s variance and sets the initial rating at 1500.
 Assume the win rate of player A is $E(\sigma^{A},\sigma^{B})$:
@@
 2.4 Performance Based RL (PBRL)
-PBRL [16] is a typical algorithm using ordinal utility in RL, which divides the entire trajectory into many sub-trajectories with fixed length $k$, and using $\sigma^{A}\succ\sigma^{B}$ indicates that human experts prefer sub-trajectory $\sigma^{A}$ to sub-trajectory $\sigma^{B}$.
-If human experts judge $\sigma^{A}\succ\sigma^{B}$, then there is a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
-$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
-So it is reasonable to assume that under the proxy reward function $\hat{r}$, the probability of an experts making $\sigma^{A}\succ\sigma^{B}$ is:
+PBRL [16] is a typical RL algorithm leveraging ordinal utility, which partitions each trajectory into fixed-length $k$ sub-trajectories, with $\sigma^{A}\succ\sigma^{B}$ denoting human experts’ preference for sub-trajectory $\sigma^{A}$ over $\sigma^{B}$.
+If experts judge $\sigma^{A}\succ\sigma^{B}$, it implies a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
+$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})$
+$>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
+Thus, under the proxy reward function $\hat{r}$, it is reasonable to assume the probability of an expert preferring $\sigma^{A}\succ\sigma^{B}$ is:
 $\hat{P}(\sigma^{A}\succ\sigma^{B})=\frac{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})}{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})+\exp\sum\limits_{i=1}^{i=k_{B}}\hat{r}(s^{B}_{i},a^{B}_{i})}$ (8)
-Under this assumption, store expert’s preference results $\left(\sigma^{A},\sigma^{B}\,\mu\right)$ into the database $\mathcal{D}$, where $\mu$ is the result function, if $\mu(A)=1$ means $\sigma^{A}\succ\sigma^{B}$, $\mu(A)=0$ means $\sigma^{B}\succ\sigma^{A}$, according to the preference, use the cross entropy loss to optimize the proxy reward function $\hat{r}$ and train reinforcement learning through the proxy reward function $\hat{r}$:
+Under this framework, expert preference results $\left(\sigma^{A},\sigma^{B},\mu\right)$ are stored in database $\mathcal{D}$, where $\mu$ denotes the preference indicator: $\mu(A)=1$ indicates $\sigma^{A}\succ\sigma^{B}$, and $\mu(A)=0$ indicates $\sigma^{B}\succ\sigma^{A}$.
+Leveraging these preferences, we optimize the proxy reward function $\hat{r}$ using cross-entropy loss and train the RL agent via the proxy reward function $\hat{r}$:
 $\mathcal{L}(\hat{r})=-\sum_{(\sigma^{A},\sigma^{B}\,\mu)\in\mathcal{D}}\left(\begin{aligned} \mu(A)\log\hat{P}(\sigma^{A}\succ\sigma^{B})+\\ \mu(B)\log\hat{P}(\sigma^{B}\succ\sigma^{A})\end{aligned}\right)$ (9)
 3 Methodology
-In this section, we introduce our new method, ELO rating based RL.
-It abandons the numerical reward, sorts the trajectories through ELO rating, and uses a simplified version of the return decomposition method to assign ELO rating to each state-action pair.
-3.1 Trail Reward Estimation Via Elo Rating
-Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose that in trajectory $\sigma^{A}$ achieves better results than trajectory $\sigma^{B}$ formulating as $\sigma^{A}\succ\sigma^{B}$ (ie.
-in Chess, player A wins Player B).
-Denote $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$ as the reward obtained by this trajectory $\sigma^{X}$, where $\overline{\mathcal{R}(\sigma^{X})}$ represents the expected reward of using the trajectory $\sigma^{X}$, $\epsilon$ denote the reward fluctuation.
-In general, it is very difficult to estimate $\overline{\mathcal{R}(\sigma^{X})}$, either need to consume huge computing resources to fully explore the entire solution space; or hire human experts to design appropriate intermediate reward to guide the agent to explore more efficiently.
-Although it is difficult to give the appropriate reward for each trajectory $\mathcal{R}(\sigma^{X})$, it is much simpler to assess the pros and cons of each trajectory.
-At this time, it is better to use the idea of ordinal utility, using $\mathcal{U}(\sigma)$ to represent the ELO rating of the trajectory $\sigma$, and use $\mathcal{U}(\sigma)$ as the reward of the entire trajectory to guide RL training.
-When $\sigma^{A}\succ\sigma^{B}$, it can be regarded as player $\sigma^{A}$ defeated player $\sigma^{B}$ ($S_{A}=1$); if the expert cannot judge the pros and cons of the two trajectories, it can be described a draw, $\sigma^{A}\approx\sigma^{B}$ ($\approx$ indicates that the experts believe that the results of the two trajectories are similar), can be regarded as a draw between the two sides ($S_{A}=0.5$) .
-If $\sigma^{A}\succ\sigma^{B}$ the event $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$ will occur with high probability, and the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ can be updated to:
+In this section, we introduce our novel method, ELO-Rating based RL (ERRL), which replaces numerical rewards with trajectory rankings via ELO ratings and employs a simplified return decomposition to assign ELO-based values to state-action pairs.
+3.1 Trajectory Reward Estimation via ELO Rating
+Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose $\sigma^{A}$ outperforms $\sigma^{B}$ (denoted $\sigma^{A}\succ\sigma^{B}$, e.g., Player A wins in chess).
+Let the trajectory reward be $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$, where $\overline{\mathcal{R}(\sigma^{X})}$ is the expected reward of trajectory $\sigma^{X}$ and $\epsilon$ models reward fluctuation.
+Estimating $\overline{\mathcal{R}(\sigma^{X})}$ is challenging: it either requires exhaustive exploration of the solution space (prohibitive in computation) or human experts to design intermediate rewards for efficient guidance.
+While assigning precise rewards $\mathcal{R}(\sigma^{X})$ is difficult, comparing trajectories is more feasible.
+Leveraging ordinal utility theory, we define $\mathcal{U}(\sigma)$ as the ELO rating of trajectory $\sigma$, using $\mathcal{U}(\sigma)$ as the trajectory reward for RL training.
+When $\sigma^{A}\succ\sigma^{B}$, this corresponds to ”player” $\sigma^{A}$ defeating ”player” $\sigma^{B}$ (with score $S_{A}=1$); if experts cannot distinguish between trajectories ($\sigma^{A}\approx\sigma^{B}$), this is treated as a draw ($S_{A}=0.5$).
+The event $\sigma^{A}\succ\sigma^{B}$ implies a high probability that $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$, allowing the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ to be updated as:
 $E(\sigma^{A},\sigma^{B})=\frac{1}{1+10^{(\hat{\mathcal{U}}(\sigma^{B})-\hat{\mathcal{U}}(\sigma^{A}))/\eta}}$ (10)
@@
 $\hat{\mathcal{U}}(\sigma^{X})=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (12)
-In value based RL, the value function itself implies a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, there is no need to redefine a proxy reward function, We use deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate long-term ELO rating on the state-action pair $(s,a)$.
-The score for the entire trajectory $\sigma^{X}$ is denote as $\hat{\mathcal{U}}(\sigma^{X})$:
+In value-based RL, the value function inherently encodes a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, obviating the need for a proxy reward function.
+We employ a deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate the long-term ELO rating for state-action pair $(s,a)$.
+The score for the entire trajectory $\sigma^{X}$ is denoted as $\hat{\mathcal{U}}(\sigma^{X})$:
 $\hat{\mathcal{U}}(\sigma^{X})$ $=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (13)
 $=\sum_{t=1}^{k_{X}}\left[\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})-\gamma\hat{G}_{\text{elo}}(s^{X}_{t+1},a^{X}_{t+1})\right]$
-The long-term reward for each state-action pair is updated as (proved in Appendix A):
+The long-term reward for each state-action pair is updated as (proved in Appendix Appendix A):
 $\hat{G}^{\prime}_{\text{elo}}(s^{X}_{t},a^{X}_{t})=\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})+\frac{K(S_{X}-E_{X})}{k_{X}}$ (14)
 3.3 Combining RL and Elo Rating
-At the code implementation level, our algorithm is compatible with mainstream RL algorithms.
-Our algorithm migrates perfectly to any RL algorithm with only a modification to the reward function.
-In Algorithm 1, we integrate actor-critic RL with Elo rating.
+At the implementation level, our algorithm is compatible with mainstream RL frameworks.
+It can be seamlessly integrated into any RL algorithm by modifying the reward function alone.
+In Algorithm 1, we demonstrate this compatibility by integrating Elo rating with an actor-critic RL framework.
 Algorithm 1 AC RL with Elo Rating
 1:Initialize $\mathcal{D}\leftarrow\varnothing$
-Predictably, the preferences of experts will have a large impact on the results of ordinal-based RL .
-In addition, the $K$ and $\eta$ parameters of ELO rating will also have a great impact on the results theoretically.
-We will discuss this problem in detail in Section 4 .
+2:for $\ell=1,2,\dots$ do
+3:     Collect a rollout trajectory $\sigma$ using the current policy, and store trajectory $\sigma$ into the replayer buffer $\mathcal{D}$.
+4:     Sample $N$ trajectories $W=\{\sigma^{i}\in\mathcal{D}\}^{N}_{i=1}$ from the replay buffer.
+5:     for $i=1,2,\dots,N$ do
+6:         Generates a shuffled integer list $Y$ from 1 to $N$.
+7:         Use equation 10 to get a new long term Elo rating estimation. $E(\sigma^{i},\sigma^{Y[i]})=1/(1+10^{(\hat{\mathcal{U}}(\sigma^{Y[i]})-\hat{\mathcal{U}}(\sigma^{i}))/\eta})$
+8:     end for
+9:     Use equation 14 to gradient update on the critic network.
+$\mathcal{L}(\hat{G}^{\prime}_{\text{elo}})=\sum_{i=1}^{N}\sum_{(s^{i},a^{i})\in\sigma^{i}}\left(\frac{K(\mu(\sigma^{i})-E(\sigma^{i},\sigma^{Y[i]}))}{k_{i}}\right)$ (15)
+10:     Use critic network to update actor network.
+11:end for
+Notably, expert preferences are expected to significantly influence the outcomes of ordinal-based RL.
+Theoretically, the ELO rating parameters $K$ and $\eta$ will also impact performance, which we will analyze in detail in Section 4.
 4 Experiments
-In this section, we conduct experiments on a suite of Atari benchmark tasks to investigate the empirical performance of our proposed method.
-We compare our method with several baseline algorithms in the literature, and conduct an ablation study.
-4.1 Performance in Atari Benchmark with episodic preference
-4.1.1 Setups
-To evaluate the behavior of ERRL, we compared its performance with several based algorithms across four Atari tasks.
-In ordinal based RL, all trajectories record some key events for expert judgment instead of returning a numerical reward.
-These tasks are long-term feedback problems (up to 5000 steps).
-We use the open source RL framework RLLib [24] as the baseline, which is an efficient parallel RL framework and adopts the standard preprocessing of Deepmind [7].
-All results have been fine-tuned, and detailed settings can be found in Appendix B.
-Since our experiment is based on ordinal RL, how to generate the ordinal model is crucial to the learning results. Here we take Pong as an example to illustrate the ordinal production process.
-Pong is a game that simulates table tennis (ping-pong).
-In normal mode, if you shot the ball out of the opponent’s baseline, you will get 1 point; if you don’t catch the opponent’s shot, you will deduct 1 point.
-The player who reaches 21 points first wins the game.
-The final trajectory reward is $\mathcal{R}(\sigma)$, its own score minus the opponent’s score.
-And our evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
+In this section, we conduct experiments on a suite of Atari benchmark tasks to evaluate the empirical performance of our proposed method.
+We compare ERRL with several state-of-the-art baselines and perform ablation studies to analyze its components.
+4.1 Experimental Setups
+To assess ERRL’s behavior, we compare its performance against multiple baseline algorithms across four Atari tasks, all featuring long-term feedback (up to 5000 steps).
+In these ordinal-based RL settings, trajectories record key events for expert preference judgments rather than returning numerical rewards.
+We use the open-source RL framework RLLib [24] as the baseline, which adopts DeepMind’s standard preprocessing pipeline [7] and enables efficient parallel training.
+All results are fine-tuned, with detailed configurations provided in Appendix Appendix B.
+As our experiments rely on ordinal RL, generating the ordinal preference model is critical for learning outcomes. We illustrate the ordinal generation process using Pong as a case study.
+[…]
   1. 1.
-    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
+    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$, then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$; otherwise, proceed to the next criterion.
   2. 2.
-    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
+    If both $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$, then $\sigma^{A}\approx\sigma^{B}$; otherwise, proceed to the third criterion.
   3. 3.
-    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
-$k$ indicates the length of the trajectory. In order to avoid the trajectory being too long, the game automatically terminates after the step reaches 5000.
-The rationale behind this design is twofold: if defeating your opponent directly is not feasible, the strategy shifts towards dragging your opponent into extended play, aiming to hit as many of the opponent’s balls as possible.
-Conversely, when you can defeat your opponent, the approach is to leverage any means necessary to secure the win.
-So there are no more restrictions.
-For more game settings please refer to Appendix B.
-4.1.2 Performance in Atari Benchmark
+    If the trajectory length $k_{A}>k_{B}$ or $k_{A}<k_{B}$, then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$; otherwise, $\sigma^{A}\approx\sigma^{B}$.
+Here, $k$ denotes the trajectory length, and the game terminates automatically after 5000 steps to prevent excessively long trajectories.
+This design follows two principles: when direct victory is unfeasible, the strategy shifts to prolonged play, aiming to maximize ball returns; when victory is achievable, any means to secure the win are prioritized.
+No additional constraints are imposed.
+For detailed game configurations, refer to Appendix Appendix B.
+4.2 Performance in Atari Benchmark
 Figure 2: Learning curves on a suite of Atari benchmark tasks with trajectory feedback.
@@
 Due to the trajectory length is indeterminate, we set up an evaluation point every 60 trajectories, about $3\times 10^{4}\sim 3\times 10^{5}$ time-steps.
-A comparison of overall performance, as shown in Figure 2, shows that ERRL generally outperforms the baseline algorithms on these several problems.
-Our method can achieve higher sample efficiency and generate better policies after convergence.
-Because the returns are sparse, it is difficult for the original PPO to learn useful information.
-As for the previous reward decomposition function, it is mainly aimed at the decomposition of numerical reward.
-The reward obtained directly from the conversion of ordinal reward will increase the fluctuation of training.
-When the agent employs PPO for training with the reward conversion based on ELO Rating, it can learn an effective policy.
-This is because the ELO rating tends to amplify the performance gap between trajectories.
-In the long-term RL problems, ELO ratings are more conducive to highlighting the pros and cons of trajectories than the compact real rewards.
-So that the agent can find useful information.
+The overall performance comparison in Figure 2 demonstrates that ERRL consistently outperforms baseline algorithms across these tasks, achieving higher sample efficiency and converging to superior policies.
+Due to the sparse nature of rewards, vanilla PPO struggles to extract meaningful information.
+Existing reward decomposition methods, designed for numerical rewards, introduce instability when applied to ordinal preferences.
+In contrast, ERRL leverages ELO ratings to convert ordinal feedback into scalable rewards, amplifying performance differences between trajectories.
+This approach proves particularly effective in long-horizon RL problems, where ELO ratings more effectively highlight trajectory quality than raw numerical rewards, enabling the agent to identify optimal strategies.
 4.3 Ablation Study
-We study the value of $\eta$ in the ERRL algorithm under the condition of a fixed learning rate of $1\times 10^{-4}$.
-For the value of $K$, we refer to the original design of ELO rating to fix $K=0.04\times\eta$.
-As described in Section 3, $\eta$ controls the fluctuation range of the ELO rating, smaller $\eta$ represents greater randomness of the zero-sum game, larger training fluctuations, but training faster; larger $\eta$ represents less randomness, and reduced training fluctuations, but training slower.
-So we need a suitable $\eta$ to balance training speed and fluctuation.
+We study the impact of $\eta$ in the ERRL algorithm under a fixed learning rate of $1\times 10^{-4}$. For the parameter $K$, we follow the original ELO design by setting $K=0.04\times\eta$.
+As discussed in Section 3, $\eta$ regulates the fluctuation range of ELO ratings: smaller $\eta$ introduces higher zero-sum game randomness, leading to larger training volatility but faster convergence; larger $\eta$ reduces randomness, decreasing training fluctuations but slowing down learning.
+Thus, an optimal $\eta$ is needed to balance training speed and stability.
 Figure 3: Learning curves of ERRL in Pong with different choices of hyper-parameter $\eta$. The curves with $\eta=0.01$ correspond to the default implementation.
-We perform an ablation study on the hyperparameter $\eta$ that represents the range of fluctuations in the ELO rating.
-Considering that the time steps of the game will change with the progress of the game, we decided to relating $\eta$ with the time-steps of the trajectory, that is, the $\eta$ actually used in the ELO Rating calculation is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$.
-For the convenience of writing, we will use $\eta$ to represent $\eta_{\text{our}}$ in the following text if there is no special emphasis.
-We evaluate ERRL with a set of choices of hyperparameter $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$.
-The experimental results shown in Figure 3 show that this hyperparameter has a great impact on RL training.
-A larger $\eta$ will cause the training fluctuation to be too large and directly collapse, while a smaller $\eta$ will cause the training to be too slow and easily lead to over fitting.
-An acceptable range of $\eta$ is in interval $[0.05,0.005]$.
-From the time-steps of the entire trajectory, regardless of the size of $\eta$, the final time-steps converge to 2100$\sim$2400.
-It can be seen that under the same preference, if the training can converge, the choice of $\eta$ does not change the final strategies, and we will discuss the relationship between strategy selection and expert preference in Appendix C.
+We conduct an ablation study on the hyperparameter $\eta$, which controls the fluctuation range of the ELO rating.
+Given that game time steps vary during gameplay, we link $\eta$ to the trajectory length, such that the effective $\eta$ value used in ELO rating calculations is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$.
+For simplicity, we use $\eta$ to denote $\eta_{\text{our}}$ throughout the remainder of the text unless otherwise specified.
+We evaluate ERRL with $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$.
+As illustrated in Figure 3, this hyperparameter significantly impacts RL training.
+Larger $\eta$ values lead to excessive training volatility and potential model collapse, whereas smaller $\eta$ values slow down convergence and increase the risk of overfitting.
+An optimal range for $\eta$ is identified as $[0.05,0.005]$.
+Across all trajectories, regardless of $\eta$’s magnitude, the final time steps converge within the range of 2100–2400.
+This indicates that, under consistent expert preferences, the choice of $\eta$ does not alter the learned optimal strategies.
+We further discuss the relationship between strategy selection and expert preferences in Appendix Appendix C.
 5 Related Work
 RL without Environmental Reward.
-While RL typically relies on environmental reward, several studies have explored RL without direct environmental rewards.
-This research can be broadly categorized into two approaches.
-The first approach focuses on capturing essential information within trajectories and using these differences to define learning rewards Eysenbach et al.
-[28] proposed a pre-training method to train the discriminator to identify which agent generated the trajectory.
-the agent always hopes that the discriminator can easily identify itself, so the agent always tends to find a unique state.
+While traditional RL relies on environmental rewards, recent studies have explored reward-free RL frameworks, which can be categorized into two main approaches.
+The first approach focuses on extracting discriminative information from trajectories to define surrogate rewards.
+Eysenbach et al.
+[28] proposed a pre-training method where a discriminator is trained to identify the agent that generated a trajectory.
+The agent aims to produce trajectories that are easily distinguishable, driving it to seek unique states.
 Hu et al.
-[26] uses multidimensional features to determine the Pareto frontier.
-Only the agents at the Pareto frontier will perform the next training.
-This approach enables game designers to create a variety of agents with distinct styles, but this type of method is generally only suitable for increasing the diversity of agents.
-Another idea is to use human embeddings in RL, such as PBRL [16] introduced in Section 2.4.
-In addition, Kimin et al.
-[13] proposed the method of PEBBLE, which has demonstrated exceptional outcomes.
-.
+[26] uses multidimensional features to determine the Pareto frontier, with only frontier agents proceeding to the next training phase.
+This approach enhances agent diversity but is typically limited to diversity-driven tasks.
+The second approach integrates human preferences into RL, such as Preference-Based RL (PBRL) [16] discussed in Section 2.4.
+Additionally, Kimin et al.
+[13] proposed PEBBLE, which has shown remarkable performance in reward-free settings.
 Sparse Rewards RL.
-Sparse rewards significantly hinder the application of RL [11, 8], with numerous studies addressing this challenge.
-A straightforward solution is the integration of human experience [1, 4].
-As long as the human benefit is designed perfectly enough, this approach can lead to successful outcomes.
-However, relying heavily on human expertise can be expensive and may inadvertently limit the capabilities of the RL agent [1, 4, 5].
-Another approach is reward decomposition, which decomposes the total reward into individual state-action pairs [15, 20].
-usually this method requires reward shaping to achieve good results.
+Sparse rewards remain a critical challenge for RL applications [11, 8], with research addressing this through multiple strategies.
+A direct solution is to incorporate human expertise [1, 4], which can yield successful results if expert knowledge is well-codified.
+However, heavy reliance on human input is costly and may inadvertently constrain agent capabilities [1, 4, 5].
+Another approach is reward decomposition, which distributes cumulative rewards to individual state-action pairs [15, 20].
+This often requires reward shaping to achieve effectiveness.
 Marcin et al.
-[10] also achieves good results by reusing historical information, but this only works for tasks with simple and clear states.
+[10] leverages historical information reuse for sparse reward tasks, though this is limited to domains with simple, well-defined states.
 6 Conclusion and Future Work
-We propose a new preference-based RL algorithm ERRL, which can achieve good performance in the 5000-step long-tern trajectory feedback RL problem.
-ERRL abandons the classical numerical reward, and estimates the ordinal reward of each trajectory through the expert’s preference and the ELO rating method.
-In addition, a new reward decomposition method is adopted to reduce the fluctuation of training in ERRL setting.
-Experiments show that ERRL outperforms previous methods in solution quality.
-In future work, we will explore the theoretical basis for combining ELO rating with RL, conduct more detailed experiments, discuss the effectiveness of our algorithm when it is difficult for experts to make judgments, and explore the feasibility of ELO rating in the remaining areas.
-Such as reward shaping  [27], games with hard-to-determine payoffs (ie.
-Texas hold’em and Hanabi) [25], or research on the diversity of RL agents [26, 28].
+We propose a novel preference-based RL algorithm, ERRL (ELO-Rating based Reinforcement Learning), which demonstrates superior performance in 5000-step long-term trajectory-feedback RL tasks.
+ERRL eschews classical numerical rewards, instead estimating ordinal rewards for each trajectory via expert preferences and the ELO rating system.
+Additionally, a new reward decomposition method is introduced to mitigate training volatility under the ERRL framework.
+Experimental results show that ERRL outperforms state-of-the-art methods in solution quality.
+For future work, we will explore the theoretical foundations of integrating ELO rating with RL, conduct more detailed ablation studies, investigate algorithm effectiveness under ambiguous expert preferences, and explore the feasibility of ELO rating in extended domains including reward shaping [27], games with complex payoffs (e.g., Texas hold’em and Hanabi) [25], and research on RL agent diversity [26, 28].
 Appendix A
@@
 According to maximum likelihood estimation, the above problem can be transformed into an optimization problem.
-The current formula  17 takes the maximum value of $x_{t}$ as the most likely update range:
+The current formula  18 takes the maximum value of $x_{t}$ as the most likely update range:
 $\begin{split}\max\prod_{t=1}^{k_{A}}\frac{e^{-\delta_{t}^{A}}}{(1+e^{-\delta_{t}^{A}})^{2}}\\ \text{s.t.}~{}~{}~{}~{}(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (18)
-Taking the logarithm of the original formula 17, there are:
+Taking the logarithm of the original formula 18, there are:
 $\max-\sum_{t=1}^{k_{A}}\left(\delta_{t}^{A}+2\ln(1+e^{-\delta_{t}^{A}})\right)$ (19)
@@
 ELO $K$ $0.04\times\eta$
-Our open source main experimental code and details can be found in https://github.com/Zealoter/ERRL, as our experiments are still ongoing, we will gradually improve our code and experimental details.
+Our open source main experimental code and details can be found in Github, as our experiments are still ongoing, we will gradually improve our code and experimental details.
 Expert Preferences